OpenStudy (anonymous):

On the moon, g=1.62 m/s . An astronaut has a weight of 600 N on the Earth. a) What is her mass on the moon ? b) what is her weight on the moon ?

OpenStudy (anonymous):

Hi Unsure how to answer this one ?

OpenStudy (anonymous):

yeah , i got for a) m = F/a = 600 N / 1.62 m/s2 = 370,4 g but im not certain if that right and how to then solve b)

OpenStudy (anonymous):

First you have to get a clear separation in your mind between earth data and moon data

OpenStudy (anonymous):

The question tells you that the astronout weighs 600N on the earth. So, if you want to use weight = mg to find the astronaut's mass, you must be consistent and use the value of g on earth with the 600N weight - do you see that ?

OpenStudy (anonymous):

The fact that you first calculated a mass of 370kg for your astronaut should set the alarm bells ringing !! That would be a seriously obese astronaut.

OpenStudy (anonymous):

So, use the weight of 600N together with earth's value of g to calculate the mass of the astronaut.

OpenStudy (anonymous):

Ok, if the mass of the astronaut on the earth is 61 kg isnt his mass the same on the moon?

OpenStudy (anonymous):

Most definitely, yes.

OpenStudy (anonymous):

So now, if you want to calculate the astronaut's weight on the moon, what formula do you need ?

OpenStudy (anonymous):

Is it W = m*g = 61 kg * 1,62 m/s2 = 98,8 N ?

OpenStudy (anonymous):

That's right, same mass everywhere, but g depends on where you are, so Wmoon = mass x gmoon

OpenStudy (anonymous):

So on the moon, objects weigh approximately one sixth of their weight on Earth

OpenStudy (anonymous):

The question tries to trick you a little bit by asking, what is her mass on the moon, because actually her mass is the same everywhere.

OpenStudy (anonymous):

thank you very much!!

OpenStudy (anonymous):

welcome